# How much has DESI dark energy evolved since DR1?

### v1

@@ Abstract
DESI has reported a dynamical dark energy (DE) signal based on the $w_{0}w_{a}$ model that is in conflict with Hubble tension. Recalling that the combination of DESI DR1 BAO and DR1 full-shape (FS) modeling are consistent with $\Lambda$CDM, in this letter we comment on the status of fluctuations in DR1 BAO documented in Adame et al. (2024a); Ó Colgáin et al. (2024) in the DR2 update. In particular, we note that neither DR1 BAO nor DR2 BAO nor DR2 BAO+CMB confronted to the $w_{0}w_{a}$ model with relaxed model parameter priors confirm late-time accelerated expansion today. Translating DESI BAO constraints into flat $\Lambda$CDM constraints, we observe that the DESI LRG1 constraint remains the most prominent outlier preferring larger $\Omega_{m}$ values, LRG2 switches from smaller to larger $\Omega_{m}$ values relative to Planck-$\Lambda$CDM, and ELG data drive the relatively low $\Omega_{m}$ in the full DR2 BAO. We observe that one cannot restore $w_{0}=-1$ within one $1\sigma$ by removing either LRG1 or LRG2, but LRG2 in DR2, in contrast to LRG1 in DR1, now has a greater bearing on $w_{0}>-1$. We conclude that the BAO has yet to stabilise, but the general trend is towards greater consistency with DESI DR1 FS modeling results, where there may be no dynamical DE signal in DESI data alone.
@@ I Introduction
Recent observations of a statistically significant dynamical dark energy (DE) Adame et al. (2024a, b); Abdul Karim et al. (2025) (see Abbott et al. (2024); Rubin et al. (2023) for earlier claims) based on the CPL model Chevallier and Polarski (2001); Linder (2003) are problematic on a number of fronts. First, local (model-independent) $H_{0}$ determinations are biased to $H_{0}>70$ km/s/Mpc values (see Skara and Perivolaropoulos (2020); Di Valentino et al. (2021); Abdalla et al. (2022); Di Valentino et al. (2025) for reviews). To the extent of our knowledge, it has been appreciated since 2018, perhaps even earlier, that a DE equation of state in the traditional quintessence regime $w_{\textrm{DE}}(z)>-1$ exacerbates the Hubble tension Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022). The result holds in simple $w_{\textrm{DE}}(z)$ parameterizations, e. g. $w$CDM Vagnozzi (2020), $w_{0}w_{a}$ Vagnozzi et al. (2018); Alestas et al. (2020), and more general field theories Banerjee et al. (2021); Lee et al. (2022). In particular, it was observed in Lee et al. (2022) that $w_{0}:=w_{\textrm{DE}}(z=0)>-1$ hinders a resolution to Hubble tension even if $w_{\textrm{DE}}(z)<-1$ at $z>0$. One sees the problem clearly in DESI DR2 results Abdul Karim et al. (2025): in the combination BAO+CMB+SNe, for different SNe samples, respectively Pantheon+ Brout et al. (2022), DES Abbott et al. (2024) and Union3 Rubin et al. (2023), as $w_{0}$ increases, $H_{0}$ decreases. This observation pertains to any combination of datasets with $w_{0}>-1$ and is more general than the CPL model, applying to DE models fitted to DESI data in lieu of CPL, e. g. Lodha et al. (2025a, b).
@@
The second problem is that given the inevitable unknown unknowns in cosmology, one can only trust a result if one sees it independently in distinct datasets. The prototypical example of this is late-time accelerated expansion, which necessitates the presence of DE modeled through $\Lambda$ in the $\Lambda$CDM model, and is seen independently across virtually all observables; not seeing support for $\Lambda$ undermines the observable. In contrast, the DESI dynamical DE claim Adame et al. (2024a, b); Abdul Karim et al. (2025) is only statistically significant when datasets are combined. In defense of combining datasets, it is well recognized that BAO, SNe and CMB only weakly constrain the CPL model on their own. This is in part due to the fact that CPL can be (or should be) viewed as a Taylor expansion in a small parameter $1-a$, which leads to inflated errors relative to other parameterizations Ó Colgáin et al. (2021). Marginalizing over higher order terms also leads to inflated errors in any $w_{0}w_{a}$ model Nesseris et al. (2025).
@@
Even when one restricts attention to independent datasets, a third problem arises: if there is a dynamical DE signal in DESI BAO Adame et al. (2024a); Abdul Karim et al. (2025), when combined with DESI full-shape (FS) modeling Adame et al. (2024b), there may be no dynamical DE signal. To appreciate this, note that Fig. 16 of Adame et al. (2024c) returns consistent constant $\Omega_{m}$ constraints with the $\Lambda$CDM model within $0.7\sigma$. Thus, it is imperative to identify BAO data points that are driving one away from the constant $\Omega_{m}$ behaviour characteristic of the $\Lambda$CDM model Ó Colgáin and Sheikh-Jabbari (2024). Note, in DESI DR2 BAO+CMB, one has a dynamical DE signal at $3.1\sigma$ Abdul Karim et al. (2025). This deviation is expected to be driven by the departures from $\Lambda$CDM behaviour in the relevant late-Universe observable, here BAO. As we shall show, DR1 BAO, DR2 BAO and DR2 BAO+CMB fail to confirm late-time accelerated expansion today.
@@
Given the obvious conflict with Hubble tension and the risk of statistical fluctuations in LRG BAO data when compared to FS modeling, we revisit earlier analysis. As we show, LRG1 continues to return a $\Lambda$CDM $\Omega_{m}$ value larger than Planck, but at $1.6\sigma$ removed, it is less anomalous. Moreover, LRG2 BAO has flipped from a lower $\Omega_{m}$ value to a higher $\Omega_{m}$ value relative to Planck. This means that while LRG1 drove the $w_{0}>-1$ result in DR1 BAO Ó Colgáin et al. (2024), the $w_{0}>-1$ in DR2 BAO is now driven by both LRG1 and LRG2, but primarily LRG2. On the other hand, whereas the full DESI DR1 BAO dataset preferred a lower $\Omega_{m}$ value relative to Planck driven by LRGs and emission line galaxies (ELGs), in DESI DR2 BAO the lower $\Omega_{m}$ value is driven exclusively by the ELGs. See Ormondroyd et al. (2025) for a recent comparison of DESI DR1 and DR2 BAO. We note that despite the failure to confirm $q_{0}<0$ (without SNe) and the risk of fluctuations, physical modeling of DR2 BAO is underway Brandenberger (2025); Luu et al. (2025); Nakagawa et al. (2025); Paliathanasis (2025); Silva et al. (2025); Wang and Piao (2025); Chaussidon et al. (2025); Kessler et al. (2025); You et al. (2025); Pan et al. (2025); Shlivko et al. (2025).
@@ II Analysis
We begin with a comment on DESI priors $w_{0}\in[-3,1],w_{a}\in[-3,2]$ Adame et al. (2024a); Abdul Karim et al. (2025) when DR2 BAO is confronted with the CPL model. Note, the choice of priors here is arbitrary as there is no theoretical guidance. In contrast, the DES collaboration use more agnostic priors $w_{0}\in[-10,5],w_{a}\in[-20,10]$ Abbott et al. (2024), despite also working with a single observable, namely SNe instead of BAO. By comparing the blue contours in Fig. 1 (DESI priors) to the blue contours in Fig. 2 (DES priors) one sees that the posteriors are skewed with narrower priors. In particular, it is worth noting that the $H_{0}r_{d}$ posterior is positively skewed, while $\Omega_{m}$ and $w_{0}$ posteriors are negatively skewed in Fig. 1. When one defines credible intervals using the most common choices, namely i) equal-tailed intervals and ii) highest density intervals with a mode central value, this skewness manifests itself in terms of larger errors coinciding with the longer tails in the posterior. In contrast, DESI appears to quote the mean as a central value Abdul Karim et al. (2025), $\Omega_{m}=0.352^{+0.041}_{-0.018}$, $w_{0}=-0.48^{+0.35}_{-0.17}$, so the smaller errors coincide with the longer tails. Replacing the mean value with the mode, we find $\Omega_{m}=0.375^{+0.020}_{-0.037}$ and $w_{0}=-0.21^{+0.08}_{-0.44}$, but otherwise we agree with the DESI $68\%$ credible intervals up to small numbers.
Figure 1: CPL model posteriors for DR2 BAO data subject to the DESI priors $w_{0}\in[-3,1]$, $w_{a}\in[-3,2]$ and $w_{0}+w_{a}<0$. The skewness in $(H_{0}r_{d},\Omega_{m},w_{0})$ posteriors comes from the $w_{a}$ bound $w_{a}\geq-3$.
@@
In Fig. 2, having relaxed the priors, we note that the posteriors are visibly more symmetric and Gaussian. Table 1 shows the corresponding parameter constraints. Given the symmetric posteriors, it makes little difference how one defines $68\%$ credible intervals, so we opt for equal-tailed intervals based on percentiles. As Fig. 2 shows, DR2 BAO data confronted to the CPL model is more consistent than DR1 BAO with late-time accelerated expansion today, which requires
@@
It is worth noting that $w_{a}$ does not appear in this expression. We examine this requirement [^1] Consistency with the 2011 Physics Nobel Prize, at least at redshift $z=0$, implies a stronger requirement: $q_{0}<0$ at a few $\sigma$ (ideally $5\sigma$ or more). DESI DR2 BAO on its own is not yet of sufficient constraining power to fulfill this requirement within the context of the CPL model. As another comment, our results here can be contrasted with other studies, e.g. Gómez-Valent (2019), where also assuming the CPL model but different data, $q_{0}<0$ is established at $2.5\sigma$. by evaluating the expression for $q_{0}$ on the MCMC chains and examine $q_{0}<0$. See Fig. 3 for the $q_{0}$ posteriors. Converting the MCMC chains into constraints on $q_{0}$, we find that DR1 BAO confronted to the CPL model rules out $q_{0}<0$ at $95.7\%$ confidence level, corresponding to $1.7\sigma$ for a one-sided Gaussian. For DR2 fitted to the CPL model, we find $q_{0}<0$ is ruled out at $76.9\%$ confidence level ($0.7\sigma$ for a one-sided Gaussian). See Wang (2024) for an earlier observation of this tension in DR1. With the improvement in data quality between DR1 and DR2, restrictive priors are no longer required and DR2 BAO shows progress in that $q_{0}$ is less positive.
Combining BAO with external datasets further alleviates this problem. However, even for CMB+DR2 BAO, one can convert $\Omega_{m}=0.353\pm 0.021$, $w_{0}=-0.42\pm 0.21$ Abdul Karim et al. (2025) into $q_{0}=0.09\pm 0.20$, so this result also fails to confirm late-time accelerated expansion today. Only when one combines DR2 BAO with SNe can one confirm late-time accelerated expansion today in a meaningful way. Note, for canonical values of $\Omega_{m}\sim 0.3$, we see that larger values of $w_{0}$ are problematic from equation (1). On the flip side, combining BAO with SNe lowers $w_{0}$ to avoid any contradiction.
@@
As an added check, we replace the CPL model with the analogous $z$-expansion model, $w_{\textrm{DE}}(z)=w_{0}+z\,w_{a}$, which may be viewed as a Taylor expansion in $z$ instead of $1-a=z/(1+z)$ for CPL. One finds a central value more consistent with late-time accelerated expansion today while all the errors shrink accordingly. This reduction in the errors is expected, since as explained in Ó Colgáin et al. (2021), $z$ is a larger expansion parameter than $1-a$ and this allows the data to place stronger constraints on $w_{a}$, which in turn better constrains the remaining parameters. Objectively, the CPL model is a dynamical DE model that is paradoxically less sensitive at lower redshifts. Changing the model from CPL to the $z$-expansion model, we find that $q_{0}<0$ is ruled out at $57.6\%$ confidence level ($0.2\sigma$ for a one-sided Gaussian). See Fig. 3.
Our next task is to identify which data points in DESI DR2 BAO are driving the $w_{0}>-1$ result. One could alternatively focus on the complementary parameter $w_{a}$. We follow the methodology of Ó Colgáin et al. (2024), where for each effective redshift in Table IV of Abdul Karim et al. (2025) with both $D_{M}(z_{i})/r_{d}$ and $D_{H}(z_{i})/r_{d}$ constraints, we construct a $2\times 2$ covariance matrix with the correlation $r$, generate 10,000 $(D_{M}(z_{i})/r_{d},D_{H}(z_{i})/r_{d})$ pairs, and solve the following equation for $\Omega_{m}$ for each pair:
@@
This ratio only depends on $\Omega_{m}$ in the $\Lambda$CDM model where $E(z)=\sqrt{1-\Omega_{m}+\Omega_{m}(1+z)^{3}}$. It has been checked that this methodology leads to comparable errors to Markov Chain Monte Carlo (MCMC) Ó Colgáin et al. (2024). The result of this exercise is shown in Table 2 and Fig. 4, where only the lower redshift bright galaxy sample (BGS) is omitted. As remarked in Abdul Karim et al. (2025), this is not a problem as we do not expect strong constraints on $\Omega_{m}$ at lower redshifts. We include the results of LRG3 and ELG1 for completeness in Table 2, but observe that they are not independent from the LRG3+ELG1 entry Abdul Karim et al. (2025).
@@
As remarked earlier, LRG1 data now leads to an $\Omega_{m}$ value that is more consistent with the traditional $\Omega_{m}\sim 0.3$. Relative to the Planck $\Omega_{m}$ value in red in Fig. 4, we see that all constraints intersect the red strip except for LRG1 and LRG3+ELG1. Splitting the LRG3+ELG1 constraint into its components, one sees that this is due to the low $\Omega_{m}$ value ELG1 tracer. Shifting the red horizontal strip downwards to the location of the green strip corresponding to the DESI DR2 BAO $\Omega_{m}$ constraint for the full sample, one can see that this provides visually a better fit to all constraints, except the LRG1 constraint, which returns a larger $\Omega_{m}$ value at $1.8\sigma$ [^2] Removing the LRG1 data we have $\Omega_{m}=0.293\pm 0.009$ in the $\Lambda$CDM model, so the discrepancy between LRG1 constraint in Table 2 and the remaining data is also $1.8\sigma$. This makes LRG1 the most prominent outlier once again. It should be noted that in contrast to DR1, where LRG2 data was contributing to the lower $\Omega_{m}$ in the full BAO sample, we can now confirm that this is driven exclusively by ELG data. It is also worth noting that the LRG3+ELG1 constraint has exhibited the smallest shift in $\Omega_{m}$ value between DR1 and DR2.
@@
We now ask again the question posed in Ó Colgáin et al. (2024): is it possible to remove a single constraint, for example LRG1, and recover $w_{0}=-1$ (or $w_{a}=0$) within $1\sigma$? In Table 3 we document the effect of removing LRG1 and LRG2 data, which are the most obvious points that could be driving the $w_{0}>-1$ result in the full sample. Interestingly, we find that the removal of LRG1 pushes one further into a regime where there is no accelerated expansion today. In contrast, we find that removing LRG2 brings the CPL parameters to within $1.3\sigma$ of $\Lambda$ ($w_{0}=-1,w_{a}=0$). What this implies is that the $w_{0}>-1$ deviation from $\Lambda$CDM in DR2 BAO is driven by LRG2 data, whereas in DR1 BAO, this was driven by LRG1 data. The reader should note that LRG1 appears to moderate the large $w_{0}$ value attributable to LRG2 data. Indeed, with LRG1 data removed, so that only LRG2 is driving the $w_{0}>-1$ deviation, we find that $q_{0}<0$ is disfavoured at $87.3\%$ confidence level ($1.1\sigma$ for a one-sided Gaussian).

### v2

@@ Abstract
DESI has reported a dynamical dark energy (DE) signal based on the $w_{0}w_{a}$ model that is in conflict with Hubble tension. Recalling that the combination of DESI DR1 BAO and DR1 full-shape (FS) modeling are consistent with $\Lambda$CDM, in this letter we comment on the status of fluctuations in DR1 BAO documented in Adame et al. (2024a); Ó Colgáin et al. (2024) in the DR2 update. In particular, we note that neither DR1 BAO nor DR2 BAO nor DR2 BAO+CMB confronted to the $w_{0}w_{a}$ model with relaxed model parameter priors confirm late-time accelerated expansion today. Translating DESI BAO constraints into flat $\Lambda$CDM constraints, we observe that the LRG1 constraint remains the most prominent outlier, a distinction now held jointly with ELG1, LRG2 switches from smaller to larger $\Omega_{m}$ values relative to Planck-$\Lambda$CDM, and ELG data drive the relatively low $\Omega_{m}$ in the full DR2 BAO. We observe that one cannot restore $w_{0}=-1$ within one $1\sigma$ by removing either LRG1 or ELG1 or LRG2, but LRG2 in DR2, in contrast to LRG1 in DR1, now has the greatest bearing on $w_{0}>-1$. We conclude that BAO has yet to stabilise, but the general trend is towards greater consistency with DESI DR1 FS modeling results, where there may be no dynamical DE signal in DESI data alone.
@@ I Introduction
Recent observations of a statistically significant dynamical dark energy (DE) Adame et al. (2024a, b); Abdul Karim et al. (2025) (see Abbott et al. (2024); Rubin et al. (2023) for earlier claims) based on the CPL model Chevallier and Polarski (2001); Linder (2003) are problematic on a number of fronts. First, local (model-independent) $H_{0}$ determinations are biased to $H_{0}>70$ km/s/Mpc values (see Skara and Perivolaropoulos (2020); Di Valentino et al. (2021); Abdalla et al. (2022); Di Valentino et al. (2025) for reviews). To the extent of our knowledge, it has been appreciated since 2018, perhaps even earlier, that a DE equation of state in the traditional quintessence regime $w_{\textrm{DE}}(z)>-1$ exacerbates the Hubble tension Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022). The result holds in simple $w_{\textrm{DE}}(z)$ parameterizations, e. g. $w$CDM Vagnozzi (2020), $w_{0}w_{a}$ Vagnozzi et al. (2018); Alestas et al. (2020), and more general field theories Banerjee et al. (2021); Lee et al. (2022). See appendix for a simple analytic argument why this must be so.
In particular, it was observed in Lee et al. (2022) that $w_{0}:=w_{\textrm{DE}}(z=0)>-1$ hinders a resolution to Hubble tension even if $w_{\textrm{DE}}(z)<-1$ at $z>0$. One sees the problem clearly in DESI DR2 results Abdul Karim et al. (2025): in the combination BAO+CMB+SNe, for different SNe samples, respectively Pantheon+ Brout et al. (2022), DES Abbott et al. (2024) and Union3 Rubin et al. (2023), as $w_{0}$ increases from $w_{0}=-1$, i. e. the $\Lambda$CDM deviation becomes more statistically significant, $H_{0}$ decreases. This observation pertains to any combination of datasets with $w_{0}>-1$ and is more general than the CPL model, applying to DE models fitted to DESI data in lieu of CPL, e. g. Lodha et al. (2025a, b).
@@
The second problem is that given the inevitable unknown unknowns in cosmology, one can only trust a result if one sees it independently in distinct datasets. The prototypical example of this is late-time accelerated expansion, which necessitates the presence of DE modeled through $\Lambda$ in the $\Lambda$CDM model, and is seen independently across virtually all observables; not seeing support for $\Lambda$ undermines the observable. In contrast, the DESI dynamical DE claim Adame et al. (2024a, b); Abdul Karim et al. (2025) is only statistically significant when datasets are combined. In defense of combining datasets, it is well recognized that BAO, SNe and CMB only weakly constrain the CPL model on their own. This is in part due to the fact that CPL can be (or should be) viewed as a Taylor expansion in a small parameter $1-a$, which leads to inflated errors relative to other parameterizations Ó Colgáin et al. (2021). Marginalizing over higher order terms also leads to inflated errors in any $w_{0}w_{a}$ model Nesseris et al. (2025). Moreover, see Roy Choudhury and Okumura (2024); Steinhardt et al. (2025) for studies highlighting uncertainty in the DESI dynamical DE claims in the extended CPL parameter space.
@@
Even when one restricts attention to independent datasets, a third problem arises: if there is a dynamical DE signal in DESI BAO Adame et al. (2024a); Abdul Karim et al. (2025), when combined with DESI full-shape (FS) modeling Adame et al. (2024b), there may be no dynamical DE signal. To appreciate this, note that Fig. 16 of Adame et al. (2024c) returns consistent constant $\Omega_{m}$ constraints with the $\Lambda$CDM model within $0.7\sigma$. Thus, it is imperative to identify BAO data points that are driving one away from the constant $\Omega_{m}$ behaviour characteristic of the $\Lambda$CDM model Ó Colgáin and Sheikh-Jabbari (2024). Note, in DESI DR2 BAO+CMB, one has a dynamical DE signal at $3.1\sigma$ Abdul Karim et al. (2025). This deviation is expected to be driven by two effects: i) departures from $\Lambda$CDM behaviour in BAO and ii) a discrepancy in the $\Lambda$CDM parameter $\Omega_{m}$ between BAO and CMB. As we shall show, DR1 BAO, DR2 BAO and DR2 BAO+CMB fail to confirm late-time accelerated expansion today.
@@
Given the obvious conflict with Hubble tension and the risk of statistical fluctuations in LRG BAO data when compared to FS modeling, we revisit earlier analysis. As we show, LRG1 continues to return a $\Lambda$CDM $\Omega_{m}$ value larger than Planck Aghanim et al. (2020), $\Omega_{m}=0.315\pm 0.07$, but at $1.6\sigma$ removed, it is less anomalous. Moreover, LRG2 BAO has flipped from a lower $\Omega_{m}$ value to a higher $\Omega_{m}$ value relative to Planck. This means that while LRG1 drove the $w_{0}>-1$ result in DR1 BAO Ó Colgáin et al. (2024), the $w_{0}>-1$ in DR2 BAO is now driven by both LRG1 and LRG2, but primarily LRG2. On the other hand, whereas the full DESI DR1 BAO dataset preferred a lower $\Omega_{m}$ value relative to Planck driven by LRGs and emission line galaxies (ELGs), in DESI DR2 BAO the lower $\Omega_{m}$ value is driven exclusively by the ELGs, in particular ELG1. See Ormondroyd et al. (2025) for a recent comparison of DESI DR1 and DR2 BAO. We note that despite the failure to confirm $q_{0}<0$ (without SNe) and the risk of fluctuations, physical modeling of DR2 BAO is underway Brandenberger (2025); Luu et al. (2025); Nakagawa et al. (2025); Paliathanasis (2025); Silva et al. (2025); Wang and Piao (2025); Chaussidon et al. (2025); Kessler et al. (2025); You et al. (2025); Pan et al. (2025); Shlivko et al. (2025).
@@ II Analysis
We begin with a comment on DESI priors $w_{0}\in[-3,1],w_{a}\in[-3,2]$ Adame et al. (2024a); Abdul Karim et al. (2025) when DR2 BAO is confronted with the CPL model. [^1] We do not consider curvature, since anomalies Handley (2021); Di Valentino et al. (2019); van Putten (2025) have been largely ameliorated by a new Planck likelihood Tristram et al. (2024). Note, the choice of priors here is arbitrary as there is no theoretical guidance. In contrast, the DES collaboration use more agnostic priors $w_{0}\in[-10,5],w_{a}\in[-20,10]$ Abbott et al. (2024), despite also working with a single observable, namely SNe instead of BAO. By comparing the blue contours in Fig. 1 (DESI priors) to the blue contours in Fig. 2 (DES priors) one sees that the posteriors are skewed with narrower priors. In particular, it is worth noting that the $H_{0}r_{d}$ posterior is positively skewed, while $\Omega_{m}$ and $w_{0}$ posteriors are negatively skewed in Fig. 1. When one defines credible intervals using the most common choices, namely i) equal-tailed intervals and ii) highest density intervals with a mode central value, this skewness manifests itself in terms of larger errors coinciding with the longer tails in the posterior. In contrast, DESI appears to quote the mean as a central value Abdul Karim et al. (2025), $\Omega_{m}=0.352^{+0.041}_{-0.018}$, $w_{0}=-0.48^{+0.35}_{-0.17}$, so the smaller errors coincide with the longer tails. Replacing the mean value with the mode, we find $\Omega_{m}=0.375^{+0.020}_{-0.037}$ and $w_{0}=-0.21^{+0.08}_{-0.44}$, but otherwise we agree with the DESI $68\%$ credible intervals up to small numbers.
Figure 1: CPL model posteriors for DR2 BAO data subject to the DESI priors $w_{0}\in[-3,1]$, $w_{a}\in[-3,2]$ and $w_{0}+w_{a}<0$. The skewness in $(H_{0}r_{d},\Omega_{m},w_{0})$ posteriors comes from the bound $w_{a}\geq-3$.
@@
In Fig. 2, having relaxed the priors, we note that the posteriors are visibly more symmetric and Gaussian. Table 1 shows the corresponding parameter constraints. Despite $(w_{0},w_{a})$ values being removed from $\Lambda$CDM expectations, assuming the Planck value $r_{d}=147.09$ Mpc to extract $H_{0}$, we find an age of the Universe $t_{U}$ for CPL fitted to DR1 and DR2 of $t_{U}=13.66\pm 0.08$ Gyrs and $t_{U}=13.69\pm 0.05$ Gyrs, repsectively. Both these values are consistent with a recent estimate from globular clusters, $t_{U}=13.57^{+0.28}_{-0.27}$ Gyrs Valcin et al. (2025). The key point is that relaxing the $(w_{0},w_{a})$ priors does not lead to a problem for the age of the Universe. Given the symmetric posteriors, it makes little difference how one defines $68\%$ credible intervals, so we opt for equal-tailed intervals based on percentiles. As Fig. 2 shows, DR2 BAO data confronted to the CPL model is more consistent than DR1 BAO with late-time accelerated expansion today, which requires
@@
It is worth noting that $w_{a}$ does not appear in this expression. We examine this requirement [^2] Consistency with the 2011 Physics Nobel Prize, at least at redshift $z=0$, implies a stronger requirement: $q_{0}<0$ at a few $\sigma$ (ideally $5\sigma$ or more). DESI DR2 BAO on its own is not yet of sufficient constraining power to fulfill this requirement within the context of the CPL model. As another comment, our results here can be contrasted with other studies, e.g. Gómez-Valent (2019), where also assuming the CPL model but different data, $q_{0}<0$ is established at $2.5\sigma$. by evaluating the expression for $q_{0}$ on the MCMC chains and check if $q_{0}<0$. See Fig. 3 for the $q_{0}$ posteriors. Converting the MCMC chains into constraints on $q_{0}$, we find that DR1 BAO confronted to the CPL model rules out $q_{0}<0$ at $95.7\%$ confidence level, corresponding to $1.7\sigma$ for a one-sided Gaussian. For DR2 BAO fitted to the CPL model, we find $q_{0}<0$ is ruled out at $76.9\%$ confidence level ($0.7\sigma$ for a one-sided Gaussian). See Wang (2024) for an earlier observation of this tension in DR1. With the improvement in data quality between DR1 and DR2, restrictive priors are no longer required and DR2 BAO shows progress in that $q_{0}$ is less positive.
Combining BAO with external datasets further alleviates this problem. However, even for CMB+DR2 BAO, one can convert $\Omega_{m}=0.353\pm 0.021$, $w_{0}=-0.42\pm 0.21$ Abdul Karim et al. (2025) into $q_{0}=0.09\pm 0.20$, so this result also fails to confirm late-time accelerated expansion today. Our analysis here assumes the $\Omega_{m}$ and $w_{0}$ parameters are uncorrelated normally distributed random variables, which can only overestimate the errors on $q_{0}$, so the discrepancy with $q_{0}<0$ is underestimated. Only when one combines DR2 BAO with SNe can one confirm late-time accelerated expansion today in a meaningful way. Note, for canonical values of $\Omega_{m}\sim 0.3$, we see that larger values of $w_{0}$ are problematic from equation (1). In contrast, combining BAO with SNe lowers $w_{0}$ to avoid any contradiction.
Interestingly, the same observation can be made for DES SNe Abbott et al. (2024), where $\Omega_{m}=0.495^{+0.033}_{-0.043}$, $w_{0}=-0.36^{+0.36}_{-0.30}$ can be converted into $q_{0}=0.23\pm 0.27$.  [^3] Whether one symmetrizes the errors by adopting the largest error as the standard deviation for a normal distribution or models the aysmmetry through distinct normal distributions for values larger and smaller than the central value, there is no appreciable difference to the outcome. Given that DESI BAO and DES SNe have a higher effective redshift than traditional SNe samples, this may highlight the difficulty constraining $q_{0}$ to the widely accepted value, $q_{0}<0$, when the datasets are poorly anchored at lower redshifts. However, theoretically, there is a more sensational possibility. The Trans-Planckian Censorship Conjecture (TCC) Bedroya and Vafa (2020); Bedroya et al. (2020) implies that late-time accelerated expansion must come to an end Brandenberger (2019, 2025); Li et al. (2025), so one could interpret $q_{0}>0$ as observational support.
As an added check, we replace the CPL model with the analogous $z$-expansion model, $w_{\textrm{DE}}(z)=w_{0}+z\,w_{a}$, which may be viewed as a Taylor expansion in $z$ instead of $1-a=z/(1+z)$ for CPL. One finds a central value more consistent with late-time accelerated expansion today while all the errors shrink accordingly. This reduction in the errors is expected, since as explained in Ó Colgáin et al. (2021), $z$ is a larger expansion parameter than $1-a$ and this allows the data to place stronger constraints on $w_{a}$, which in turn better constrains the remaining parameters. Objectively, the CPL model is a dynamical DE model that is paradoxically less sensitive at lower redshifts. Changing the model from CPL to the $z$-expansion model, we find that $q_{0}<0$ is ruled out at $57.6\%$ confidence level ($0.2\sigma$ for a one-sided Gaussian). See Fig. 3 for the posterior.
@@
Our next task is to identify which data points in DESI DR2 BAO are driving the $w_{0}>-1$ result. One could alternatively focus on the complementary parameter $w_{a}$. The logic we use is that deviations from $\Lambda$CDM captured by the $w_{0}w_{a}$ model must be evident as differences in $\Omega_{m}$ at different redshifts in the flat $\Lambda$CDM model. We follow the methodology of Ó Colgáin et al. (2024), where for each effective redshift in Table IV of Abdul Karim et al. (2025) with both $D_{M}(z_{i})/r_{d}$ and $D_{H}(z_{i})/r_{d}$ constraints, we construct a $2\times 2$ covariance matrix with the correlation $r$, generate 10,000 $(D_{M}(z_{i})/r_{d},D_{H}(z_{i})/r_{d})$ pairs, and solve the following equation for $\Omega_{m}$ for each pair:
@@
This ratio only depends on $\Omega_{m}$ in the $\Lambda$CDM model where $E(z)=\sqrt{1-\Omega_{m}+\Omega_{m}(1+z)^{3}}$. It has been checked that this methodology leads to comparable errors to Markov Chain Monte Carlo (MCMC) Ó Colgáin et al. (2024). The result of this exercise is shown in Table 2 and Fig. 4, where only the lower redshift bright galaxy sample (BGS) is omitted. For each entry in Table 2 we confirmed that MCMC gives comparable results, as expected. As remarked in Abdul Karim et al. (2025), this is not a problem as we do not expect strong constraints on $\Omega_{m}$ at lower redshifts. We include the results of LRG3 and ELG1 for completeness in Table 2, but observe that they are not independent from the LRG3+ELG1 entry Abdul Karim et al. (2025).
@@
As remarked earlier, LRG1 data now leads to an $\Omega_{m}$ value that is more consistent with the traditional $\Omega_{m}\sim 0.3$. Relative to the Planck $\Omega_{m}$ value in red in Fig. 4, we see that all constraints intersect the red strip except for LRG1 and LRG3+ELG1. Splitting the LRG3+ELG1 constraint into its components, one sees that this is due to the low $\Omega_{m}$ value ELG1 tracer. Shifting the red horizontal strip downwards to the location of the green strip corresponding to the DESI DR2 BAO $\Omega_{m}$ constraint for the full sample, one can see that this provides visually a better fit to all constraints, except the LRG1 constraint and ELG1 constraint once separated from LRG3. LRG1 returns a larger $\Omega_{m}$ value at $1.8\sigma$ and ELG1 returns a smaller $\Omega_{m}$ value also at $1.8\sigma$ [^4] More carefully, removing the LRG1 and ELG1 data from the remaining data we have $\Omega_{m}=0.293\pm 0.009$ and $\Omega_{m}=0.301\pm 0.009$, respectively, so the discrepancies with the LRG1 and ELG1 constraints in Table 2 and the remaining data are $1.8\sigma$ and $1.9\sigma$. This makes LRG1 the most prominent outlier once again, but ELG1, when separated from LRG3, is on par. It should be noted that in contrast to DR1, where LRG2 data was contributing to the lower $\Omega_{m}$ in the full BAO sample, we can now confirm that this is driven exclusively by ELG data. Thus, if one wanted to restore consistency with Planck-$\Lambda$CDM, it would be enough to explore ELG tracers, in particular ELG1, for potential systematics. That being said, while DR2 BAO prefers an $\Omega_{m}$ value $1.6\sigma$ lower than Planck, removing ELG1 leads to an $\Omega_{m}$ value that is $1.2\sigma$ lower. This can be compared with DR1 BAO, which led to an $\Omega_{m}$ value $1.2\sigma$ lower than Planck. It is also worth noting that the LRG3+ELG1 constraint has exhibited the smallest shift in $\Omega_{m}$ value between DR1 and DR2.
We now ask again the question posed in Ó Colgáin et al. (2024): is it possible to remove a single constraint, for example LRG1, and recover $w_{0}=-1$ (or $w_{a}=0$) within $1\sigma$? In Table 3 we document the effect of removing LRG1, LRG2 and ELG1 data, which are the most obvious outliers that could contribute to the $w_{0}>-1$ result in the full sample. From the CPL DR2 BAO entry in Table 1 we observe that $w_{0}=-1$ is $1.9\sigma$ from the $w_{0}$ value preferred by the full sample. From Table 3, removing LRG1, ELG1 and LRG2 in turn, we find $w_{0}=-1$ is $1.9\sigma$, $1.7\sigma$ and $1.3\sigma$ removed, respectively.
@@
Interestingly, we find that the removal of LRG1 pushes one further into a regime where there is no accelerated expansion today, but has no effect on the preference for $w_{0}>-1$. That being said, we find that $q_{0}<0$ is disfavoured at $87.3\%$ confidence level ($1.1\sigma$ for a one-sided Gaussian). Removing ELG1 brings one marginally closer to $w_{0}=-1$, confirming that this outlier has some influence on the $w_{0}>-1$ result. In contrast, we find that removing LRG2 brings $w_{0}$ closest to $w_{0}=-1$. What this implies is that the $w_{0}>-1$ deviation from $\Lambda$CDM in DR2 BAO is driven mainly by LRG2 data, whereas in DR1 BAO, this was driven by LRG1 data.